On June 30, 2017, Wisconsin, Inc., issued $137,250 in debt and 22,300 new shares of its $10 par value stock to Badger Company owners in exchange for all of the outstanding shares of that company. Wisconsin shares had a fair value of $40 per share. Prior to the combination, the financial statements for Wisconsin and Badger for the six-month period ending June 30, 2017, were as follows:

Wisconsin Badger
Revenues $ (983,000) $ (477,000)
Expenses 673,000 293,000
Net income $ (310,000) $ (184,000)
Retained earnings, 1/1 $ (849,000) $ (238,000)
Net income (310,000) (184,000)
Dividends declared 92,500 0
Retained earnings, 6/30 $ (1,066,500) $ (422,000)
Cash $ 170,500 $ 165,000
Receivables and inventory 452,000 211,000
Patented technology (net) 917,000 398,000
Equipment (net) 745,000 617,000
Total assets $ 2,284,500 $ 1,391,000
Liabilities $ (588,000) $ (499,000)
Common stock (360,000) (200,000)
Additional paid-in capital (270,000) (270,000)
Retained earnings (1,066,500) (422,000)
Total liabilities and equities $ (2,284,500) $ (1,391,000)


Wisconsin also paid $32,000 to a broker for arranging the transaction. In addition, Wisconsin paid $42,600 in stock issuance costs. Badger’s equipment was actually worth $728,750, but its patented technology was valued at only $371,900.


What are the consolidated balances for the following accounts?

1. Goodwill?

2. Common Stock?

3. Additional paid-in capital?
Answer #1

Part 1

Consideration transferred (fair value) for Badger Stock (137250+(22300*40))

1029250

Book Value of Badger, 6/30 (200000+270000+422000)

892000

Fair Value in Excess of Book Value

137250

Excess fair value (undervalued equipment) (728750-617000)

111750

Excess fair value (overvalued patented technology) (371900-398000)

(26100)

Goodwill (137250-111750+26100)

$51600

Part 2

Common stock = parent's book value after recording the newly‑issued shares = 360000+(22300*10) = $583000

Part 3

Additional paid-in capital = parent's book value after recording the newly‑issued shares = 270000+(22300*30)-42600 = $896400